OpenStudy (anonymous):

Find the exact area of the surface obtained by rotating the curve about the x-axis. y=sqrt(1+3x), 1

OpenStudy (anonymous):

Hi I can help u

OpenStudy (anonymous):

the 2nd radical isn't quite right for surface area integral, multiply by dA dA=1+f′(x)2−−−−−−−−√ f′(x)=24x+1−−−−−√ →1+44x+1−−−−−−−−−√=4x+54x+1−−−−−−√

OpenStudy (anonymous):

your welcome...for an answer i got 98pi/3

OpenStudy (anonymous):

medal and fan me plz

OpenStudy (anonymous):

98pi/3 isn't correct

OpenStudy (thadyoung):

First, draw the graph, then set up your integral. Here is what you have: